OpenStudy (anonymous):

Trying to understand a mixed arithmetic operations with fractions problem, need explanation.

OpenStudy (anonymous):

Evaluate \[\frac{ 5 }{ 8} \times \frac{ 1 }{ 3 } + \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

6/74

OpenStudy (anonymous):

I don't want the answer. I'm trying to understand the process of finding out the answer. I don't understand how to solve it. @mrs.mathteachergifygi

OpenStudy (anonymous):

what eves

OpenStudy (anonymous):

You're so helpful, thanks... >.>